# Reproducibility and Artifact Consistency of the SIGIR 2022 Recommender Systems Papers Based on Message Passing

### v1

@@ Abstract.
Graph-based techniques relying on neural networks and embeddings have gained attention as a way to develop Recommender Systems (RS) with several papers on the topic presented at SIGIR 2022 and 2023. Given the importance of ensuring that published research is methodologically sound and reproducible, in this paper we analyze 10 graph-based RS papers, most of which were published at SIGIR 2022, and assess their impact on subsequent work published in SIGIR 2023. Our analysis reveals several critical points that require attention: (i) the prevalence of bad practices, such as erroneous data splits or information leakage between training and testing data, which call into question the validity of the results; (ii) frequent inconsistencies between the provided artifacts (source code and data) and their descriptions in the paper, causing uncertainty about what is actually being evaluated; and (iii) the preference for new or complex baselines that are weaker compared to simpler ones, creating the impression of continuous improvement even when, particularly for the Amazon-Book dataset, the state-of-the-art has significantly worsened. Due to these issues, we are unable to confirm the claims made in most of the papers we examined and attempted to reproduce.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Manuscript Under Review; 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: General and reference Evaluation
@@ 1. Introduction
Defining what reproducibility means for IR and RS is still an open issue. However, despite the terminological debate, it is generally understood as the ability to obtain results very similar to those reported in the paper, whether in terms of the absolute values of the effectiveness scores or the actual labels (or recommendation lists) produced by the model (Breuer et al., 2020a; Maistro et al., 2023; Breuer et al., 2020b). One of the first issues that arise is ensuring that the original source code and data artifacts provided by the authors are, in fact, consistent with what was described in the paper. There can be several sources of inconsistencies, and indeed we report several examples of this, from erroneous data splits (an issue already identified by Ferrari Dacrema et al. (2021)) to changes in the model implementation after the publication of the paper. These types of inconsistencies are problematic because they hide what is actually being reproduced, whether it is what is described in the paper or something else, and may easily propagate to subsequent publications if other researchers use those artifacts themselves. Once the consistency of the artifacts has been established, the experimental evaluation can proceed to obtain the results and determine whether they reproduce what is reported in the original paper according to the desired criteria. This approach focuses on the reproduced method but does not consider the context in which the result was presented. Indeed, in machine learning research the results are generally interpreted in a relative manner; that is, the absolute value of an effectiveness metric does not provide much information on its own but becomes meaningful when compared with the same effectiveness score measured for other methods. Typically, this is used to support claims that the proposed method is able to outperform the state-of-the-art in a given scenario. This opens a new challenge that goes beyond reproducing the absolute results, that is, confirming the conclusions that the proposed method is indeed competitive with the state-of-the-art. This requires to reevaluate several aspects of the experimental protocol, from the training procedure used for the proposed method to the way in which the state-of-the-art baselines were selected and optimized. Several previous studies over the years have shown that a large number of papers present methods that are, in fact, not competitive against simple baselines due to various bad practices, such as poor baseline optimization, information leakage, and anomalous data splits  (Armstrong et al., 2009; Ferrari Dacrema et al., 2019, 2021; Rendle et al., 2020; Kharazmi et al., 2016; Lv et al., 2021; Yang et al., 2019; Anelli et al., 2021; Ferrari Dacrema et al., 2020). Indeed, back in 2009  Armstrong et al. (2009) raised the issue by observing how new methods in IR were not competitive against some older and simpler baselines. More than 10 years later, this was found to still be the case by Kharazmi et al. (2016) and Lv et al. (2021). Similar issues have been reported in the RS domain by Ferrari Dacrema et al. (2019, 2021), who observed that it was not possible to reproduce several deep learning techniques in the RS domain and that most of them were not competitive with a set of simple fine-tuned baselines when applied to the traditional top-n collaborative filtering task.
@@
  • Are the experimental methodologies used in the original papers, as well as the publicly available artifacts, correct and consistent? While 90% of the papers provided publicly available artifacts, we identified several major inconsistencies related to anomalous data splitting. The consistency of model implementations is generally good, with rare exceptions. However, the early-stopping process is often inconsistent with the descriptions in the papers, with some cases even using test data to select the number of training epochs.
  • Can we confirm that the proposed methods are competitive against simple but robust baselines, beyond those reported in the original papers? We observe that most methods fail to outperform simple baselines, particularly on the Amazon-Book dataset, where message-passing models fall substantially behind. In some cases, the analyzed method remains highly competitive, being outperformed by only a small margin by a single baseline. However, we also find several instances where a simple ItemKNN largely outperforms them.
@@ 2.1. Terminology
While over the years there has been a lot of discussion, which is still ongoing, on how reproducibility and replicability should be defined (De Roure, 2014; Plesser, 2018), we follow the definitions by the updated ACM Policy on Artifact and Review Badging, [^1] https://www.acm.org/publications/policies/artifact-review-and-badging-current which are aligned with those of the International Vocabulary for Metrology (VIM) (Joint Committee for Guides in Metrology (2008), JCGM) and the NISO definitions for reproducibility badging (NISO, 2021). These are also the definitions adopted by the ACM SIGIR Artifact Badging committee: [^2] https://sigir.org/general-information/acm-sigir-artifact-badging/
@@
We did not perform only reproducibility; since we also verified the source code, its availability, documentation, as well as its consistency with the paper, we conducted, in the terms of the ACM SIGIR Artifact Badging committee:
@@
Note that, as explained earlier, different papers used different datasets or the same datasets but with different preprocessing or splitting. As a consequence, most of the methods examined can not be directly compared to determine which performs better and under which conditions, even when the experiment uses the same dataset. In order to address this issue, we also conduct an experiment similar to what done in (Anelli et al., 2023), where we independently optimize all the hyperparameters of the methods we attempt to reproduce on a limited set of datasets. Although such an analysis on all the originally reported datasets or on a more ample set of datasets and conditions would be appropriate and interesting, it focuses on the generalizability of the results and therefore falls outside the scope of this study and is left for future work. Moreover, such an analysis would be extremely computationally expensive. However, since we fine-tuned our baselines for each of the datasets used, they still provide a reference to qualitatively assess the relative merits of the methods examined.
@@ 2.2. Selection of Candidate Papers
Reproduciblity Study on SIGIR 2022 Papers
@@ 2.3. Consistency and Reproducibility Analysis
Regarding the first requirement, an important aspect to consider is whether reproducing a paper also requires to reproduce possible methodological mistakes that were made in the original implementation, e.g., information leakage. In those instances, we argue that running experiments which contain errors has little scientific value, therefore we correct errors when present and describe them in our analysis. This issue typically arises when selecting the optimal number of epochs to train a machine learning model. Often published papers report the optimal number of epochs but do not explain how it was obtained or, in some cases, the provided implementation performs early-stopping on the test data.
@@ 2.4. Baselines
  • NegHOSLIM (EN): linear full-rank model similar to SLIM that includes higher-order interactions as input-features (Steck and Liang, 2021). [^9] Due to the large memory requirement we trained it by using an ElasticNet loss (EN) instead of the originally proposed one, in a similar way as SLIM.
@@ 3. Detailed Analysis
The reported results are the product of extensive experiments. We report the results of approximately 800 trained models, which required to fit approximately 25.000 models during the hyperparameter optimization phase. Overall the experiments required a total computation time of 4 years. Due to the extensive number of experiments and datasets, we report only one table for each paper selecting the dataset in which we obtained the worse outcome in terms of the reproducibility of the results. Notice that this does not necessarily mean worse results, but rather the results that are furthest from those published. [^11] There are two exceptions to this: (i) Movielens 100k for GDE, because the dataset is very small, and since there are no other datasets in which we could successfully run the experiment but not reproduce the results, we select Gowalla, which is the dataset with the highest number of interactions and successfully reproduced results; (ii) Last-FM for HAKG because we had to apply a significant alteration to the dataset, and therefore we select the dataset with the second worse results in terms of reproducibility. To complement the main paper, we also provide an online appendix with extensive additional information: the full results for each dataset with up to 26 baseline algorithms, the description and statistics of the datasets as well as a table listing the hyperparameter values of the analyzed methods that we used in our experiments, specifying for each where they were described (i.e., paper or source code). [^12] Will be provided on GitHub in case of acceptance https://drive.google.com/drive/folders/12i9pohws2RpBe60XDer0np8ErHrTxI55?usp=share_link Each table groups baselines into two categories: those with closed-form solutions and those that require iterative training.
@@ 3.2. Less is More: Reweighting Important Spectral Graph Features for Recommendation
During our experiments we observed that GDE is numerically unstable, frequently failing to train properly, particularly on certain datasets. We believe this instability is related to an issue in how the lowest eigenvalues are computed. Consider a matrix containing the user-item interactions defined as $R\in\mathbb{R}^{u\times i}$, where $u$ is the number of users and $i$ is the number of items. GDE requires to compute both the highest and the lowest eigenvalues, along with their corresponding eigenvectors, for the item-item similarity matrix $S_{I-I}=R^{T}R$ and the user-user similarity matrix $S_{U-U}=RR^{T}$. While computing the highest eigenvalues can be done rather efficiently, computing the lowest ones for such large matrices is a much more computationally intensive task and is subject to numerical rounding errors. The original implementation uses the LOBPCG method (Knyazev, 2001) [^16] The implementation is based on the lobpcg function from PyTorch, see the reference documentation here https://pytorch.org/docs/stable/generated/torch.lobpcg.html which is very fast but assumes the input matrix to be positive definite, i.e., all its eigenvalues are strictly positive. Crucially, this requirement is never verified and, as we will show, it is not met. Any matrix $A$ that can be represented as $A=B^{T}B$ is at least positive semi-definite. However, to be strictly positive definite an additional condition must be met, matrix $B$ must have linearly independent columns (Strang, 2014, Ch. 7, p. 396). The number of linearly independent columns and rows of a matrix is its rank, which can never be larger than the smallest of its dimensions, i.e., $rank(R)\leq min(u,i)$. This immediately implies that at least one of the two similarity matrices will not be positive definite, simply because $R$ is rectangular. In a typical scenario $u>i$ hence $S_{U-U}$ will not be positive definite. To validate our statements, we compute the rank of the matrix $R$ containing the training set by applying the SVD method, i.e., the rank of the matrix is equal to the number of non-zero singular values. [^17] We applied the function matrix_rank from PyTorch, see the reference documentation here https://pytorch.org/docs/stable/generated/torch.linalg.matrix_rank.html The results are that on Movielens 100k and CiteULike-a there are fewer than 10 linearly dependent rows or columns, while for Movielens 1M and Pinterest there are approximately 500 and for Gowalla almost 900.181818 The detailed results are the following: Movielens 100k (users=943, items=1682, rank=940), Movielens 1M (users=6040, items=3952, rank=3401), CiteULike-a (users=5551, items=16980, rank=5544), Pinterest (users=37501, items=9836, rank=9303) and Gowalla (users=29858, items=40981, rank=28965).
@@
We also identified other methodological issues. In the original data splits, a small number of interactions (11 for CiteULike-a, which is 0.02% of the training set, and 7884 for Pinterest, which is 3.94% of the training set) appear in both the training and test sets. In our experiments we removed these interactions from the training set to avoid any overlap and prevent the possibility of information leakage. Furthermore, although the source code reports the optimal number of training epochs for each dataset, the paper does not explain how these numbers were determined. The original implementation does not apply early-stopping, but rather trains for a large number of epochs and periodically evaluates the model on the test data. To address this, we conducted two experiments: one using the reported number of epochs and the other applying our own early-stopping methodology. We also observe that, due to the data splitting methodology, only 1% of the data is used for validation, which could result in noisy hyperparameter tuning and early-stopping. Lastly, the paper does not mention that the source code employs a definition of Recall where the denominator is not the number of relevant items for the user but rather the minimum between this value and the length of the recommendation list. While this definition has been used in previous studies (Liang et al., 2018), and is intended to normalize the metric in cases where the number of relevant items exceeds the length of the recommendation list, the existence of two competing definitions of a metric poses obstacles to reproducibility when the paper does not explicitly state which definition is being used. For our experiments, we use both the normalized definition of Recall used in the GDE paper to enable a direct comparison, as well as the more common non normalized definition of Recall.
@@
Our early-stopping runs of GDE show competitive results only on the two Movielens datasets compared to our baselines. On the other hand, our runs of GDE with the number of epochs from the paper are competitive on all datasets except CiteULike-a and Pinterest, where convergence is not achieved and the normalized Recall is around 0.002. Given the unstable nature of GDE, and its fast training time, we decided to perform a new hyperparameter optimization. With our new set of hyperparameters, GDE was able to reach convergence on Pinterest and outperform our baselines as well, with a normalized Recall of 0.1082 compared to the 0.1067 reached by the best performing baseline iALS. Finally, the results reported in the paper for GDE are competitive with our baselines across all datasets. The results obtained with a non normalized implementation of Recall are consistent with those obtained with the normalized definition used in the original source code.

### v2

@@ Abstract.
Graph-based techniques relying on neural networks and embeddings have gained attention as a way to develop Recommender Systems (RS) with several papers on the topic presented at SIGIR 2022 and 2023. Given the importance of ensuring that published research is methodologically sound and reproducible, in this paper we analyze 10 graph-based RS papers, most of which were published at SIGIR 2022, and assess their impact on subsequent work published in SIGIR 2023. Our analysis reveals several critical points that require attention: (i) the prevalence of bad practices, such as erroneous data splits or information leakage between training and testing data, which call into question the validity of the results; (ii) frequent inconsistencies between the provided artifacts (source code and data) and their descriptions in the paper, causing uncertainty about what is actually being evaluated; and (iii) the preference for new or complex baselines that are weaker compared to simpler ones, creating the impression of continuous improvement even when, particularly for the Amazon-Book dataset, the state-of-the-art has significantly worsened. Due to these issues, we are unable to confirm the claims made in most of the papers that we examined and attempted to reproduce.
@@
copyright: acmcopyrightjournal: TOISjournalyear: 2025journalvolume: 44journalnumber: 1article: 23publicationmonth: 12doi: 10.1145/3772275ccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: General and reference Evaluation
@@ 1. Introduction
Defining what reproducibility means for IR and RS is still an open issue. However, despite the terminological debate, it is generally understood as the ability to obtain results very similar to those reported in the paper, whether in terms of the absolute values of the effectiveness scores or the actual labels (or recommendation lists) produced by the model (Breuer et al., 2020a; Maistro et al., 2023; Breuer et al., 2020b). One of the first issues that arises is ensuring that the original source code and data artifacts provided by the authors are, in fact, consistent with what was described in the paper. There can be several sources of inconsistencies, and indeed we report several examples of this, from erroneous data splits (an issue already identified by Ferrari Dacrema et al. (2021)) to changes in the model implementation after the publication of the paper. These types of inconsistencies are problematic because they hide what is actually being reproduced, whether it is what is described in the paper or something else, and may easily propagate to subsequent publications if other researchers use those artifacts themselves. Once the consistency of the artifacts has been established, the experimental evaluation can proceed to obtain the results and determine whether they reproduce what is reported in the original paper according to the desired criteria. This approach focuses on the reproduced method but does not consider the context in which the result was presented. Indeed, in machine learning research the results are generally interpreted in a relative manner; that is, the absolute value of an effectiveness metric does not provide much information on its own but becomes meaningful when compared with the same effectiveness score measured for other methods. Typically, this is used to support claims that the proposed method is able to outperform the state-of-the-art in a given scenario. This opens a new challenge that goes beyond reproducing the absolute results, that is, confirming the conclusions that the proposed method is indeed competitive with the state-of-the-art. This requires reevaluating several aspects of the experimental protocol, from the training procedure used for the proposed method to the way in which the state-of-the-art baselines were selected and optimized. Several previous studies over the years have shown that a large number of papers present methods that are, in fact, not competitive against simple baselines due to various bad practices, such as poor baseline optimization, information leakage, and anomalous data splits  (Armstrong et al., 2009; Ferrari Dacrema et al., 2019, 2021; Rendle et al., 2020; Kharazmi et al., 2016; Lv et al., 2021; Yang et al., 2019; Anelli et al., 2021; Ferrari Dacrema et al., 2020; Shehzad et al., 2025; Benigni et al., 2025). Indeed, back in 2009  Armstrong et al. (2009) raised the issue by observing how new methods in IR were not competitive against some older and simpler baselines. More than 10 years later, this was found to still be the case by Kharazmi et al. (2016) and Lv et al. (2021). Similar issues have been reported in the RS domain by Ferrari Dacrema et al. (2019, 2021), who observed that it was not possible to reproduce several deep learning techniques in the RS domain and that most of them were not competitive with a set of simple fine-tuned baselines when applied to the traditional top-n collaborative filtering task.
@@
  • Are the experimental methodologies used in the original papers, as well as the publicly available artifacts, correct and consistent? While 90% of the papers provided publicly available artifacts, we identified several major inconsistencies related to anomalous data splitting. The consistency of model implementations is generally good, with rare exceptions. However, the early-stopping process is often inconsistent with the descriptions in the papers, and in some cases even relies on test data to select the number of training epochs.
  • Can we confirm that the proposed methods are competitive against simple yet robust baselines, beyond those reported in the original papers? We observe that most methods fail to outperform simple baselines, particularly on the Amazon-Book dataset, where message-passing models fall substantially behind. In some cases, the analyzed method remains highly competitive, being outperformed by only a small margin by a single baseline. However, we also find several instances where a simple ItemKNN largely outperforms them.
@@ 2.1. Terminology
While over the years there has been a lot of discussion, which is still ongoing, on how reproducibility and replicability should be defined (De Roure, 2014; Plesser, 2018), we follow the definitions by the updated ACM Policy on Artifact and Review Badging, [^1] https://www.acm.org/publications/policies/artifact-review-and-badging-current which are aligned to those of the International Vocabulary for Metrology (VIM) (Joint Committee for Guides in Metrology (2008), JCGM) and the NISO definitions for reproducibility badging (NISO, 2021). These are also the definitions adopted by the ACM SIGIR Artifact Badging committee: [^2] https://sigir.org/general-information/acm-sigir-artifact-badging/
@@
We did not perform only a reproducibility analysis; since we also verified the source code, its availability, documentation, as well as its consistency with the paper, we conducted, in the terms of the ACM SIGIR Artifact Badging committee:
@@
Note that, as explained earlier, different papers used different datasets or the same datasets but with different preprocessing or splitting. As a consequence, most of the methods examined can not be directly compared to determine which is more effective and under which conditions, even when the experiment uses the same dataset. In order to address this issue, we also conduct an experiment similar to what done in (Anelli et al., 2023), where we independently optimize all the hyperparameters of the methods we attempt to reproduce on a limited set of datasets. Although such an analysis on all the originally reported datasets or on a more ample set of datasets and conditions would be appropriate and interesting, it focuses on the generalizability of the results and therefore falls outside the scope of this study and is left for future work. Moreover, such an analysis would be extremely computationally expensive. However, since we fine-tuned our baselines for each of the datasets used, they still provide a reference to qualitatively assess the relative merits of the methods examined.
@@ 2.2. Selection of Candidate Papers
Reproducibility Study on the SIGIR 2022 Papers
@@ 2.3. Consistency and Reproducibility Analysis
Regarding the first requirement, an important aspect to consider is whether reproducing a paper also requires us to reproduce possible methodological mistakes that were made in the original implementation, e.g., information leakage. In those instances, we argue that running experiments which contain errors has little scientific value, therefore we correct errors when present and describe them in our analysis. This issue typically arises when selecting the optimal number of epochs to train a machine learning model. Often published papers report the optimal number of epochs but do not explain how it was obtained or, in some cases, the provided implementation performs early-stopping on the test data.
@@ 2.4. Baselines
  • NegHOSLIM (EN): linear full-rank model similar to SLIM, which includes higher-order interactions as input-features (Steck and Liang, 2021). [^9] Due to the large memory requirement we trained it by using an ElasticNet loss (EN) instead of the originally proposed one, in a similar way as SLIM.
@@ 3. Detailed Analysis
The reported results are the product of extensive experiments. We report the results of approximately 800 trained models, which required the fitting of approximately 25.000 models during the hyperparameter optimization phase. Overall the experiments required a total computation time of 4 years. Due to the extensive number of experiments and datasets, we report only one table for each paper selecting the dataset in which we obtained the worse outcome in terms of the reproducibility of the results. Notice that this does not necessarily mean worse results, but rather the results that are furthest from those published. [^11] There are two exceptions to this: (i) MovieLens 100k for GDE, because the dataset is very small, and since there are no other datasets in which we could successfully run the experiment but not reproduce the results, we select Gowalla, which is the dataset with the highest number of interactions and successfully reproduced results; (ii) Last-FM for HAKG because we had to apply a significant alteration to the dataset, and therefore we select the dataset with the second worst results in terms of reproducibility. To complement the main paper, we also provide extensive additional information in the appendix: the full results for each dataset with up to 26 baseline algorithms, the description and statistics of the datasets as well as a table listing the hyperparameter values of the analyzed methods that we used in our experiments, specifying for each where they were described (i.e., paper or source code). [^12] https://github.com/remaplab/TOIS25_Reproducibility-SIGIR22-GCN Each table groups baselines into two categories: those with closed-form solutions and those that require iterative training.
@@ 3.2. Less is More: Reweighting Important Spectral Graph Features for Recommendation
During our experiments we observed that GDE is numerically unstable, frequently failing to train properly, particularly on certain datasets. We believe this instability is related to an issue in how the lowest eigenvalues are computed. Consider a matrix containing the user-item interactions defined as $R\in\mathbb{R}^{u\times i}$, where $u$ is the number of users and $i$ is the number of items. GDE requires computing both the highest and the lowest eigenvalues, along with their corresponding eigenvectors, for the item-item similarity matrix $S_{I-I}=R^{T}R$ and the user-user similarity matrix $S_{U-U}=RR^{T}$. While computing the highest eigenvalues can be done rather efficiently, computing the lowest ones for such large matrices is a much more computationally intensive task and is subject to numerical rounding errors. The original implementation uses the LOBPCG method (Knyazev, 2001) [^16] The implementation is based on the lobpcg function from PyTorch, see the reference documentation here https://pytorch.org/docs/stable/generated/torch.lobpcg.html which is very fast but assumes the input matrix to be positive definite, i.e., all its eigenvalues are strictly positive. Crucially, this requirement is never verified and, as we will show, it is not met. Any matrix $A$ that can be represented as $A=B^{T}B$ is at least positive semi-definite. However, to be strictly positive definite an additional condition must be met, matrix $B$ must have linearly independent columns (Strang, 2014, Ch. 7, p. 396). The number of linearly independent columns and rows of a matrix is its rank, which can never be larger than the smallest of its dimensions, i.e., $rank(R)\leq min(u,i)$. This immediately implies that at least one of the two similarity matrices will not be positive definite, simply because $R$ is rectangular. In a typical scenario $u>i$ hence $S_{U-U}$ will not be positive definite. To validate our statements, we compute the rank of the matrix $R$ containing the training set by applying the SVD method, i.e., the rank of the matrix is equal to the number of non-zero singular values. [^17] We applied the function matrix_rank from PyTorch, see the reference documentation here https://pytorch.org/docs/stable/generated/torch.linalg.matrix_rank.html The results are that on MovieLens 100k and CiteULike-a there are fewer than 10 linearly dependent rows or columns, while for MovieLens 1M and Pinterest there are approximately 500 and for Gowalla almost 900.181818The detailed results are the following: MovieLens 100k (users=943, items=1682, rank=940), MovieLens 1M (users=6040, items=3952, rank=3401), CiteULike-a (users=5551, items=16980, rank=5544), Pinterest (users=37501, items=9836, rank=9303) and Gowalla (users=29858, items=40981, rank=28965).
@@
We also identified other methodological issues. In the original data splits, a small number of interactions (11 for CiteULike-a, which is 0.02% of the training set, and 7884 for Pinterest, which is 3.94% of the training set) appear in both the training and test sets. In our experiments we removed these interactions from the training set to avoid any overlap and prevent the possibility of information leakage. Furthermore, although the source code reports the optimal number of training epochs for each dataset, the paper does not explain how these numbers were determined. The original implementation does not apply early-stopping, but rather trains for a large number of epochs and periodically evaluates the model on the test data. To address this, we conducted two experiments: one using the reported number of epochs and the other applying our own early-stopping methodology. We also observe that, due to the data splitting methodology, only 1% of the data is used for validation, which could result in noisy hyperparameter tuning and early-stopping. Lastly, the paper does not mention that the source code employs a definition of Recall where the denominator is not the number of relevant items for the user but rather the minimum between this value and the length of the recommendation list. While this definition has been used in previous studies (Liang et al., 2018), and is intended to normalize the metric in cases where the number of relevant items exceeds the length of the recommendation list, the existence of two competing definitions of a metric poses obstacles to reproducibility when the paper does not explicitly state which definition is being used. For our experiments, we use both the normalized definition of Recall used in the GDE paper to enable a direct comparison, as well as the more common non-normalized definition of Recall.
@@
Our early-stopping runs of GDE show competitive results only on the two MovieLens datasets compared to our baselines. On the other hand, our runs of GDE with the number of epochs from the paper are competitive on all datasets except CiteULike-a and Pinterest, where convergence is not achieved and the normalized Recall is around 0.002. Given the unstable nature of GDE, and its fast training time, we decided to perform a new hyperparameter optimization. With our new set of hyperparameters, GDE was able to reach convergence on Pinterest and outperform our baselines as well, with a normalized Recall of 0.1082 compared to the 0.1067 reached by the best performing baseline iALS. Finally, the results reported in the paper for GDE are competitive with our baselines across all datasets. The results obtained with a non-normalized implementation of Recall are consistent with those obtained with the normalized definition used in the original source code.